# Ambipolar doping of a charge-transfer insulator in the Emery model

### v1

@@ VII Redistribution of the orbital character of the dopants
Figure 8: Partial occupation of the copper $3d$ orbital, $n_{d}$ [panel (a)], and of the equivalent oxygen $2p_{x},2p_{y}$ orbitals, $n_{p}$ [panel (b)], vs doping $\delta$ for two values of temperature above and below the pseudogap-metal critical endpoint. Left: electron doping, right: hole doping. Data are for the set of parameters of Eq. 19 and $U_{d}=11.1$.
@@
First, let us consider zero doping. In the absence of the hybridization, $t_{pd}=0$, the interaction strength $U_{d}$ splits the copper $3d$ level into two levels at $\epsilon_{d}$ and $\epsilon_{d}+U_{d}$. In a charge-transfer insulator the copper level $\epsilon_{d}+U_{d}$ lies above the oxygen $2p$ band, hence the copper occupation $n_{d}$ is equal to $1$ and the oxygen occupation $n_{p}$ is equal to $2$. Hence the total occupation is equal to $n_{\rm tot}=n_{d}+2n_{p}=5$. Due to the finite hybridization, $t_{pd}\neq 0$, the charge-transfer insulator is then realised for $n_{d}=1+\eta$ and correspondingly $n_{p}=2-\eta$ (for the set of parameters in Fig. 8 one finds $n_{d}\approx 1.274$ and $n_{p}\approx 1.863$). In other words, the magnitude of the partial occupation is slightly more than half filling for $n_{d}$ and is almost fully filled for $n_{p}$, as schematically represented in the central panel of Fig. 1(b).
@@ VIII.1 $U_{d}-\delta$ phase diagram
Figure 11: (a) $U_{d}$ vs $\mu$ phase diagram at $T=1/40$ and for the set of parameters of Eq. 19. Charge-transfer insulator (CTI) phase is shown in dark blue, bounded by the dashed dark blue line with open squares. Emerging from this phase is the pseudogap (PG) phase shown in red, bounded by a dashed, red line with open red triangles. The correlated metal (CM) phase is shown in orange bounded by a dotted orange line with open orange triangles. Since the temperature of the critical end point decreases with increasing $U_{d}$, our scan of the $U_{d}-T-\mu$ space at constant temperature either cross the coexistence region (hatched) or the Widom line (open purple circles). (b) $U_{d}$ vs $\delta$ phase diagram at $T=1/40$. The CTI area of panel (a) collapses into a line at zero doping. The coexistence region of panel (a) maps into a region of coexistence (hatched with colored background) or into a forbidden region (hatched with white background). Left: electron doping, right: hole doping.
@@
To summarise, Fig. 11 indicates that the $U_{d}-T-\delta$ diagram hosts a first order transition surface. This transition surface progressively extends from zero doping to finite doping upon increasing $U_{d}$. This transition terminates in a second order critical line from which a surface of supercritical crossover emerges. This finding extends to electron doping the results of Ref. [44] which focused on hole doping only.
The phase diagram in Fig. 11 enable us to gain a further understanding of the microscopic mechanism shaping the doping driven metal to charge-transfer insulator transition. First, the $U_{d}-T-\delta$ phase diagram does not show a qualitative asymmetry upon electron or hole doping. A first order transition crosses $\delta=0$ plane and extends either doping sides. Hence, this first order transition is a robust feature against electron and hole dopings. A particle-hole symmetry only arises in the temperature and doping position of this transition. The pseudogap to metal transition at either particle and hole doping (see $T-\delta$ phase diagram for $U_{d}=11.1$ in Fig. 5) are indeed connected in the $U_{d}-T-\delta$ plane, indicating their common origin in the Mott physics in two dimensions. Second, Fig. 11 shows that the pseudogap phase emerges only above a threshold value of $U_{d}$ for opening a charge-transfer insulator at $\delta=0$, i.e. for $U>U_{d,{\rm MIT}}$, upon either electron or hole doping. This again suggests the origin of the pseudogap in the Mott physics in two dimensions, and that the pseudogap is a robust feature against electron or hole doping.
VIII.2 Local density of states
@@ VIII.3 Redistribution of the orbital character of the dopants
We can identify three trends in the slopes of the curves. (i) Below $U_{d,\rm MIT}$, the slope of the curves is large and positive, for either electron or hole doping. (ii) Above $U_{d,\rm MIT}$ and for hole doping, the slope of the curves slightly increases with increasing $U_{d}$. Equivalently, the curves have large and positive slope for small oxygen hole content, and a small and almost flat slope for higher oxygen hole content. (iii) Correspondingly, above $U_{d,\rm MIT}$ and for electron doping, the slope of the curves becomes slightly less steep with increasing $U_{d}$. Equivalently, the curves have larger slopes for small oxygen hole content, and almost a vertical slope for high oxygen hole content.
@@ IX Discussion
Our results shows that the mere introduction of holes or electrons into an ‘ionic’ charge-transfer insulator does produce an asymmetry in the doping evolution of the orbital character of the dopants, but it is not sufficient to produce a qualitative asymmetry in the nature of the doping driven transition in the normal state. Hence, we are led to conclude that by merely doping with holes or electrons a charge-transfer insulator, one cannot take into account the marked electron-hole asymmetry observed in the normal state $T-\delta$ phase diagram of real cuprates.
Our results rule out that the source of the observed electron-hole asymmetry lies solely in doping with holes or electrons the same parent charge-transfer insulator, but cannot identify the source of that asymmetry. Thus, let us discuss separately the main insights that we can derive from our work about the normal state of hole-doped cuprates, electron-doped cuprates, and the cuprates for which ambipolar doping (i.e., doping with either holes or electrons) is experimentally achievable.
@@ IX.1 Hole-doped cuprates
Similarly to the normal state of hole-doped cuprates, our results shows a charge-transfer insulator at $\delta=0$, followed by a pseudogap region [99] whose mechanism is associated to Mott physics plus short-range correlations [4, 100, 101, 102, 87]. The pseudogap found in our calculations shows a striking particle-hole asymmetric shape with respect to the Fermi energy (Figs. 6(d), 7(b), 12(d), 13(b)), as observed in experiments [103]. The pseudogap to metal transition is followed at high doping by a maximum of thermodynamic entropy vs doping, whose position is almost doping independent and extends up to high temperature (Fig. 5), as found in experiments [78, 79]. Using the same methodology of the present work, Ref. [44] has further shown that, for hole doping, the pseudogap to metal transition is actually concealed by a superconducting dome, and that the onset temperature of the pseudogap, $T^{*}$, ends abruptly is in the middle of the superconducting region, in agreement with experimental results [104, 105].
@@
Our present work also shows that doped holes mainly enter the oxygen orbitals (Fig. 9), as found in experiments [59, 60, 61, 62, 63, 64, 65, 66]. It also shows that upon increasing $U_{d}$, the copper hole content increases and the oxygen hole content decreases (Fig. 14(a)). It is interesting to compare our results in Fig. 14(a) to the experimental findings of Refs. [65, 93] for different families of cuprates. First, the parent compound of hole-doped cuprates has a ratio of copper/oxygen hole content ranging approximately from $/0.5$ 0.5 of a Tl-based cuprate family [65, 93] to $0.82/0.18$ of LSCO [65, 93]. Hence, our results in Fig. 14(a) at large value of $U_{d}$ seem to be more appropriate to describe the phenomenology realised by LSCO. Second, prior work [48, 94] has identified an anticorrelation between the oxygen hole content and the size of charge-transfer gap (i.e., the larger the oxygen hole content, the smaller the size of the charge-transfer gap). Since $U_{d}$ scales with the size of the gap, our results capture this trend (i.e., the smaller the $U_{d}$, and hence the smaller the charge-transfer gap, the higher the oxygen hole content). Third, the experimental trends for the slope of the curves close to the $\delta=0$ line and for hole-doped materials are opposite to our findings: experimentally, the curves have a small slope for small oxygen hole content (e.g. LSCO [65, 93]), and a large slope for higher oxygen content (e.g. Tl-based cuprate family [65, 93]). This suggests that to describe this experimental trend, one cannot just simply reduce $U_{d}$ in the Emery model, but other effects, such as those produced by tuning the bare charge-transfer energy $|\epsilon_{p}-\epsilon_{d}|$ and the hybridization $t_{pd}$, should be considered. We leave this investigation for further work.
@@ IX.2 Electron-doped cuprates
Let us now turn to the electron-doped case. It is long known that the phenomenology of real electron-doped materials such as PCCO and NCCO differs from that of real hole-doped materials. Let us stress that our works focuses on the normal state CDMFT solution of the Emery model, i.e. we do not allow symmetry breaking states. Physically, this means that the normal state occurs at all temperatures. In sharp contrast to experiments, our work finds a qualitative symmetry in the nature of electron and hole doped side of the $T-\delta$ normal state phase diagram (see Fig. 5). Hence we are led to conclude that the charge-transfer insulator regime studied in our theory work is not the relevant regime for the parent compound of actual electron-doped cuprate materials such as PCCO and NCCO. This also implies that the strongly correlated pseudogap arising from Mott physics in 2D found in our work is not relevant for the the pseudogap found in electron-doped cuprates such as PCCO and NCCO [106, 107], which is associated not to Mott physics but to a fluctuating precursors of a long-range antiferromagnetic order [108].
@@
Second, the high oxygen hole content in electron-doped materials correlates with their smaller charge-transfer gap at zero doping [65, 109]. This is compatible with prior work suggesting that electron-doped materials are less correlated than their hole-doped counterpart [55, 69, 8, 17]. Our finding that $U_{d}$ is one possible microscopic mechanism governing the redistribution of electrons and holes is compatible with this viewpoint. This is because (i) below $U_{d,{\rm MIT}}$ we found no strongly correlated pseudogap (see Fig. 11) and (ii) because decreasing $U_{d}$ leads to an increase of the oxygen hole content (see Fig. 14(a)). In that respect, our findings are compatible with studies [8, 17] indicating that actual electron-doped cuprates may lie on the metallic side of the metal to charge-transfer insulator boundary of the Zaanen-Sawatsky-Allen diagram. Furthermore, studies [32] on the Emery model set on the metallic side of the metal to charge-transfer insulator boundary show a better electronic screening and hence less electronic correlations with increasing electron doping. Note that our work focuses only on the effects of $U_{d}$. However other effects, such as those produced by tuning the bare charge-transfer energy $|\epsilon_{p}-\epsilon_{d}|$, can change the effective interaction strength, hence further work is needed to clarify the interplay of these effects. We leave this for further investigations.
@@ IX.3 The ambipolar Sr1-xLaxCuO2+y cuprate films
First, using scanning transmission electron microscopy (STEM), Ref. [9] shows that the high-energy features of the SLCO bandstructure remains essentially symmetric upon either electron or hole doping, despite the location of the doped carriers (doped holes mainly enter into the oxygens, doped electrons mainly enter into the copper). In particular, the charge-transfer gap remains essentially constant (around $1.3$ eV), regardless the hole vs electron nature of the carriers. This echoes our findings that (i) the distance between the charge-transfer band and the upper Hubbard band does not vary much with doping (see Fig. 7(d), 13(d)) and (ii) ambipolar doping of a well developed (i.e. an ‘ionic’) charge-transfer gap, no qualitative asymmetry arises in the nature of the doping driven transition in the normal state (see Fig. 5). Prior mean-field calculations [16] also found an intriguing electron-hole symmetry in the bandstructure of a doped charge-transfer insulator, irrespective of the location of the doped carriers.
Second, our findings suggest that doping of an ‘ionic’ charge-transfer gap should give rise to a strongly correlated pseudogap – i.e. a pseudogap associated purely with Mott physics in 2D – upon either electron or hole doping (see Fig. 5). Intriguingly, the STEM spectra of SLCO cuprate films also display pseudogap features upon either electron or hole doping (see supplemental material of Ref. [9]). Furthermore, and similar to our results (see Figs. 6, 7(a,b), 12, 13(a,b)), the STEM spectra show a pseudogap with a two-peak structure characterised by a marked asymmetry about the Fermi energy, for either electron or hole doping. Both in experiments and in our results, the electron-hole asymmetry of the pseudogap around the Fermi energy depends on the nature of the dopant carriers: for hole doping the pseudogap has a more pronounced peak below the Fermi energy, whereas for electron doping the pseudogap has a more pronounced peak above the Fermi energy. Therefore, upon changing between electron and hole doping, the electron-hole asymmetry of the pseudogap is reversed (see Figs. 7(c), 13(c)). Hence our work provides a possible theoretical framework to explain the experimental results of Ref. [9].

### v2

@@ VII Redistribution of the orbital character of the dopants
Figure 9: Partial occupation of the copper $3d$ orbital, $n_{d}$ [panel (a)], and of the equivalent oxygen $2p_{x},2p_{y}$ orbitals, $n_{p}$ [panel (b)], vs doping $\delta$ for two values of temperature above and below the pseudogap-metal critical endpoint. Left: electron doping, right: hole doping. Inset in each panel zooms into the pseudogap to metal transition region, for $T=1/60$. Data are for the set of parameters of Eq. 19 and $U_{d}=11.1$.
@@
First, let us consider zero doping. In the absence of the hybridization, $t_{pd}=0$, the interaction strength $U_{d}$ splits the copper $3d$ level into two levels at $\epsilon_{d}$ and $\epsilon_{d}+U_{d}$. In a charge-transfer insulator the copper level $\epsilon_{d}+U_{d}$ lies above the oxygen $2p$ band, hence the copper occupation $n_{d}$ is equal to $1$ and the oxygen occupation $n_{p}$ is equal to $2$. Hence the total occupation is equal to $n_{\rm tot}=n_{d}+2n_{p}=5$. Due to the finite hybridization, $t_{pd}\neq 0$, the charge-transfer insulator is then realised for $n_{d}=1+2\eta$ and correspondingly $n_{p}=2-\eta$ (for the set of parameters in Fig. 9 one finds $n_{d}\approx 1.274$ and $n_{p}\approx 1.863$). In other words, the magnitude of the partial occupation is slightly more than half filling for $n_{d}$ and is almost fully filled for $n_{p}$, as schematically represented in the central panel of Fig. 1(b).
@@ VIII.1 $U_{d}-\delta$ phase diagram
Figure 12: (a) $U_{d}$ vs $\mu$ phase diagram at $T=1/40$ and for the set of parameters of Eq. 19. Gray filled circles indicate the datapoints studied in our work. Charge-transfer insulator (CTI) phase is shown in dark blue, bounded by the dashed dark blue line with open squares. Emerging from this phase is the pseudogap (PG) phase shown in red, bounded by a dashed, red line with open red triangles. The correlated metal (CM) phase is shown in orange bounded by a dotted orange line with open orange triangles. Since the temperature of the critical end point decreases with increasing $U_{d}$, our scan of the $U_{d}-T-\mu$ space at constant temperature either cross the coexistence region (hatched) or the Widom line (open purple circles). (b) $U_{d}$ vs $\delta$ phase diagram at $T=1/40$. The CTI area of panel (a) collapses into a line at zero doping. The coexistence region of panel (a) maps into a region of coexistence (hatched with colored background) or into a forbidden region (hatched with white background). Left: electron doping, right: hole doping.
@@
To summarise, Fig. 12 indicates that the $U_{d}-T-\delta$ diagram hosts a surface of first order transitions. This transition surface progressively extends from zero doping to finite doping upon increasing $U_{d}$. This transition terminates in a second order critical line from which a surface of supercritical crossover emerges. This finding extends to electron doping the results of Ref. [44] which focused on hole doping only.
The phase diagram in Fig. 12 enable us to gain a further understanding of the microscopic mechanism shaping the doping-driven metal to charge-transfer insulator transition. First, the $U_{d}-T-\delta$ phase diagram does not show a qualitative asymmetry upon electron or hole doping. A first order transition crosses the $\delta=0$ plane and extends on either doping sides. Hence, this first order transition is a robust feature against electron and hole dopings. A particle-hole symmetry only arises in the temperature and doping position of this transition. The pseudogap to metal transition at either particle and hole doping (see $T-\delta$ phase diagram for $U_{d}=11.1$ in Fig. 5) are indeed connected in the $U_{d}-T-\delta$ plane, indicating their common origin in the Mott physics in two dimensions. Second, Fig. 12 shows that the pseudogap phase emerges only above a threshold value of $U_{d}$ for opening a charge-transfer insulator at $\delta=0$, i.e. for $U>U_{d,{\rm MIT}}$, upon either electron or hole doping. This again suggests the origin of the pseudogap in the Mott physics in two dimensions, and that the pseudogap is a robust feature against electron or hole doping.
VIII.2 Local density of states and spectral function at the Fermi level
@@
Fig. 15 shows the intensity plots of the spectral functions $A_{d}({\bf k})$ and $A_{p}({\bf k})$ at the Fermi level at different values of doping, for $U_{d}=12$ and $T=1/40$. Overall, the spectral intensity is smaller at $U_{d}=12$ than at $U_{d}=11.1$, as expected from the increased correlation strength. As discussed in Fig. 8, the pseudogap is characterized by a strong ${\bf k}$ dependent spectral functions [first row], with spectral intensity concentrated along the zone boundary at electron doping and along the zone diagonal at hole doping. Note however, that the features of the spectral intensity evolve with doping more gradually in Fig. 15 than in Fig. 8: this is because in Fig. 15 the spectral intensity is shown at temperatures above the critical endpoint of the pseudogap to metal transition.
@@ VIII.3 Redistribution of the orbital character of the dopants
We can identify three trends in the slopes of the curves close the $\delta=0$ line. (i) Below $U_{d,\rm MIT}$, the slope of the curves is large and positive, for either electron or hole doping. (ii) Above $U_{d,\rm MIT}$ and for hole doping, the slope of the curves slightly increases with increasing $U_{d}$. Equivalently, the curves have large and positive slope for small oxygen hole content, and a small and almost flat slope for higher oxygen hole content. (iii) Correspondingly, above $U_{d,\rm MIT}$ and for electron doping, the slope of the curves becomes slightly less steep with increasing $U_{d}$. Equivalently, the curves have larger slopes for small oxygen hole content, and almost a vertical slope for high oxygen hole content.
@@ IX Discussion
Our results show that the mere introduction of holes or electrons into an ‘ionic’ charge-transfer insulator does produce an asymmetry in the doping evolution of the orbital character of the dopants, but it is not sufficient to produce a qualitative asymmetry in the nature of the doping-driven transition in the normal state. Hence, we are led to conclude that by merely doping with holes or electrons a charge-transfer insulator, one cannot take into account the marked electron-hole asymmetry observed in the normal state $T-\delta$ phase diagram of real cuprates.
Our results rule out that the source of the observed electron-hole asymmetry lies solely in doping with holes or electrons the same parent charge-transfer insulator, but they cannot identify the source of that asymmetry. Thus, let us discuss separately the main insights that we can derive from our work about the normal state of hole-doped cuprates, electron-doped cuprates, and the cuprates for which ambipolar doping (i.e., doping with either holes or electrons) is experimentally achievable.
@@ IX.1 Hole-doped cuprates
Similarly to the normal state of hole-doped cuprates, our results shows a charge-transfer insulator at $\delta=0$, followed by a pseudogap region [117] whose mechanism is associated to Mott physics plus short-range correlations [4, 103, 118, 119, 92]. The pseudogap found in our calculations shows a striking particle-hole asymmetric shape with respect to the Fermi energy (Figs. 6(d), 7(b), 13(d), 14(b)), as observed in experiments [120]. The pseudogap shows a strongly momentum dependent spectral intensity (Figs. 8(a,b,i,j), 15(a,b,i,j)), with weight concentrated along the zone diagonal, in favourable agreement with the Fermi arcs revealed by angle-resolved photoemission (ARPES) experiments [121, 122]. The pseudogap to metal transition is followed at high doping by a maximum of thermodynamic entropy vs doping, whose position is almost doping independent and extends up to high temperature (Fig. 5), as found in experiments [78, 79]. Using the same methodology as the present work, Ref. [44] has further shown that, for hole doping, the pseudogap to metal transition is actually concealed by a superconducting dome, and that the onset temperature of the pseudogap, $T^{*}$, ends abruptly is in the middle of the superconducting region, in agreement with experimental results [123, 124].
@@
Our present work also shows that doped holes mainly enter the oxygen orbitals (Fig. 10), as found in experiments [59, 60, 61, 62, 63, 64, 65, 66]. It also shows that upon increasing $U_{d}$, the copper hole content increases and the oxygen hole content decreases (Fig. 16(a)). It is interesting to compare our results in Fig. 16(a) to the experimental findings of Refs. [65, 111] for different families of cuprates. First, the parent compound of hole-doped cuprates has a ratio of copper/oxygen hole content ranging approximately from $/0.5$ 0.5 of a Tl-based cuprate family [65, 111] to $0.82/0.18$ of LSCO [65, 111]. Hence, our results in Fig. 16(a) at large value of $U_{d}$ seem to be more appropriate to describe the phenomenology realised by LSCO. Second, prior work [48, 112] has identified an anticorrelation between the oxygen hole content and the size of charge-transfer gap (i.e., the larger the oxygen hole content, the smaller the size of the charge-transfer gap). Since $U_{d}$ scales with the size of the gap, our results capture this trend (i.e., the smaller the $U_{d}$, and hence the smaller the charge-transfer gap, the higher the oxygen hole content). Third, the experimental trends for the slope of the curves close to the $\delta=0$ line and for hole-doped materials are opposite to our findings: experimentally, the curves have a small slope for small oxygen hole content (e.g. LSCO [65, 111]), and a large slope for higher oxygen content (e.g. Tl-based cuprate family [65, 111]). This suggests that to describe this experimental trend, one cannot just simply reduce $U_{d}$ in the Emery model, but other effects, such as those produced by tuning the bare charge-transfer energy $|\epsilon_{p}-\epsilon_{d}|$, the hybridization $t_{pd}$, as well as the interaction strength $U_{p}$ on the oxygens should be considered. We leave this investigation for further work.
@@ IX.2 Electron-doped cuprates
Let us now turn to the electron-doped case. It is long known that the phenomenology of real electron-doped materials such as PCCO and NCCO differs from that of real hole-doped materials. Let us stress that our works focuses on the normal state CDMFT solution of the Emery model, i.e. we do not allow symmetry breaking states. Physically, this means that the normal state occurs at all temperatures. In sharp contrast to experiments, our work finds a qualitative symmetry in the nature of electron and hole doped side of the $T-\delta$ normal state phase diagram (see Fig. 5). Hence we are led to conclude that the charge-transfer insulator regime studied in our theory work is not the relevant regime for the parent compound of actual electron-doped cuprate materials such as PCCO and NCCO. This also implies that the strongly correlated pseudogap arising from Mott physics in 2D found in our work is not relevant for the the pseudogap found in electron-doped cuprates such as PCCO and NCCO [125, 126], which is associated not to Mott physics but to a fluctuating precursors of a long-range antiferromagnetic order [127]. For example, the doping evolution of the spectral function found in our calculations is not captured by ARPES experiments on NCCO [121, 122].
@@
Second, the high oxygen hole content in electron-doped materials correlates with their smaller charge-transfer gap at zero doping [65, 128]. This is compatible with prior work suggesting that electron-doped materials are less correlated than their hole-doped counterpart [55, 69, 8, 17]. Our finding that $U_{d}$ is one possible microscopic mechanism governing the redistribution of electrons and holes is compatible with this viewpoint. This is because (i) below $U_{d,{\rm MIT}}$ we found no strongly correlated pseudogap (see Fig. 12) and (ii) because decreasing $U_{d}$ leads to an increase of the oxygen hole content (see Fig. 16(a)). In that respect, our findings are compatible with studies [8, 17, 69] indicating that actual electron-doped cuprates may lie on the metallic side of the metal to charge-transfer insulator boundary of the Zaanen-Sawatsky-Allen diagram. Furthermore, studies [32] on the Emery model set on the metallic side of the metal to charge-transfer insulator boundary show a better electronic screening and hence less electronic correlations with increasing electron doping. Note that our work focuses only on the effects of $U_{d}$. However other effects, such as those produced by tuning the bare charge-transfer energy $|\epsilon_{p}-\epsilon_{d}|$, can change the effective interaction strength, hence further work is needed to clarify the interplay of these effects. We leave this for further investigations.
Finally, it is interesting to mention that an experimental advance in the annealing method, known as “protect annealing”, has shown the possibility to suppress antiferromagnetic fluctuations and thus the antiferromagnetic pseudogap in the electron doped Pr1.3-xLa0.7CexCuO4 (PLCCO) single crystals [129, 130]. The suppression of fluctuating precursors of long range antiferromagnetic order could reduce the asymmetry between the temperature vs electron doping and temperature vs hole doping phase diagram of cuprates [130]. Therefore the CDMFT normal state results could be relevant for protected annealed PCCO, as already pointed out in Ref. 108 in the context of single band 2D Hubbard model. Furthermore, Refs. [129, 131, 130] pointed out a possible discrepancy between the nominal and actual doping concentrations in electron-doped cuprates: using the Luttinger number as the doping axis, one obtains a $T-\delta$ phase diagram of electron-doped cuprates, in which the doping range of the antiferromagnetic and superconducting phases are similar to that of hole-doped cuprates.
@@ IX.3 The ambipolar Sr1-xLaxCuO2+y cuprate films
First, using scanning transmission electron microscopy (STEM), Ref. [9] shows that the high-energy features of the SLCO bandstructure remains essentially symmetric upon either electron or hole doping, despite the location of the doped carriers (doped holes mainly enter into the oxygens, doped electrons mainly enter into the copper). In particular, the charge-transfer gap remains essentially constant (around $1.3$ eV), regardless the hole vs electron nature of the carriers. This echoes our findings that (i) the distance between the charge-transfer band and the upper Hubbard band does not vary much with doping (see Fig. 7(d), 14(d)) and (ii) ambipolar doping of a well developed (i.e. an ‘ionic’) charge-transfer gap, no qualitative asymmetry arises in the nature of the doping-driven transition in the normal state (see Fig. 5). Prior mean-field calculations [16] also found an intriguing electron-hole symmetry in the bandstructure of a doped charge-transfer insulator, irrespective of the location of the doped carriers.
Second, our findings suggest that doping of an ‘ionic’ charge-transfer gap should give rise to a strongly correlated pseudogap – i.e. a pseudogap associated purely with Mott physics in 2D – upon either electron or hole doping (see Fig. 5). Intriguingly, the STEM spectra of SLCO cuprate films also display pseudogap features upon either electron or hole doping (see supplemental material of Ref. [9]). Furthermore, and similar to our results (see Figs. 6, 7(a,b), 13, 14(a,b)), the STEM spectra show a pseudogap with a two-peak structure characterised by a marked asymmetry about the Fermi energy, for either electron or hole doping. Both in experiments and in our results, the electron-hole asymmetry of the pseudogap around the Fermi energy depends on the nature of the dopant carriers: for hole doping the pseudogap has a more pronounced peak below the Fermi energy, whereas for electron doping the pseudogap has a more pronounced peak above the Fermi energy. Therefore, upon changing between electron and hole doping, the electron-hole asymmetry of the pseudogap is reversed (see Figs. 7(c), 14(c)). Hence our work provides a possible theoretical framework to explain the experimental results of Ref. [9].